# A coin flip game and generalizations of Fibonacci numbers

### v1

@@ 1. Introduction
While all the above research work and other related online discussions occurred in 2024, Simonson discussed in a book published in 2022 [6, Ch. 2] a somewhat similar (but different) game in which a carnie offers \$5 to you while in a carnival and asks you to pay \$$n$ back, where $n$ is the number of times you flip a coin until you have two consecutive heads. Let $E$ be the expected value of coin flips to end this game. If the first flip gives a tail, then the expected value of flips to get two heads in a row will be $1+E$. If the first flip gives a heads and the next is a tails, then the expected value of flips will be $2+E$. If the first two flips both give a heads then the expected flips is $2$. It follows that $E=\frac{1}{2}(1+E)+\frac{1}{4}(2+E)+\frac{2}{4}$. This implies $E=6$, so in average you will lose \$1 each time when you play this game.
There is another way to compute $E$. Let $E_{n}$ be the number of outcomes from flipping a coin $n$ times such that the game will end right after $n$ flips. This turns out to be the Fibonacci number $F_{n-1}$ since $E_{n}=n-1$ for $n=1,2$ and for $n\geq 3$, if the first flip is a heads, then the next flip must be a tails, so by induction, $E_{n}=F_{n-3}+F_{n-2}=F_{n-1}$. It follows that the expected value of $n$ is $\sum_{n=1}^{\infty}nF_{n-1}/2^{n}$, which must equal $6$. This can also be proved by standard generating function technique, although Simonson did not address that, perhaps due to the scope of his book [6].
Simonson [6, Ch. 2] also mentioned that, since you will lose money eventually by playing the first game a number of times, what if you are offered to switch roles with the carnie but using HT instead of HH to end the game? Similarly as the above, one can show that the expected number of flips to end this game is $E=4$, so you should not get tricked again by the carnie. Interestingly, as mentioned by Simonson [6, Ch. 2], although the expected number of coin flips to get HH is more than the expected number of coin flips to get HT (and thus it is more likely for HT to outnumber HH in a long run), it is equally likely to get HT or HH when either occurs, since it is a $1/2$ chance to get either a heads or a tails after a heads.
@@ 3. Generalizations of Fibonacci numbers
Recall that the Fibonacci number of order $k$ is defined by $F_{n}^{k}=0$ for $n<k-1$, $F_{n}^{k}=1$ for $n=k-1$, and $F_{n}^{k}=F_{n-1}^{k}+F_{n-2}^{k}+\cdots+F_{n-k}^{k}$ for $n\geq k$. We have $F_{n,1}=1$ for $n\geq 0$, $F_{n,2}=F_{n}$ is the well-known Fibonacci number [7, A000045], $F_{n,3}$ is the tribonacci number [7, A000073], $F_{n,4}$ is the tetranacci number [7, A000078], and so on.
@@
Special cases include $\overline{F}_{n}^{1}=n+1$, $\overline{F}_{n}^{2}=F_{n+2}-1$ [7, A000071], $\overline{F}_{n}^{3}$ [7, A008937], $\overline{F}_{n}^{4}$ [7, A107066], and so on.
@@
For $m=0$ we have $F_{n}^{k,0}=2F_{n-1}^{k,0}-F_{n-k-1}^{k,0}$, so $F_{n}^{k,0}=\overline{F}_{n}^{k}$. The case $k=m=1$ [7, A005314] and the case $(k,m)=(2,1)$ [7, A059633] have also been studied before from other perspectives.
@@
We have $\widetilde{F}_{n}^{0,0}=F_{n+2}$, $\widetilde{F}_{n}^{k,0}=F_{n+1}^{k,1}$, and some other interesting special cases of $\widetilde{F}_{n}^{k,m}$, such as $(k,m)=(0,1)$ [7, A006053], $(k,m)=(0,2)$ [7, A158943], and $(k,m)=(1,1)$ [7, A112575].
@@
Special cases include $G_{n}^{2}=\widetilde{F}_{n-1}^{1,1}$ [7, A112575] and $G_{n}^{3}=F_{n-2}^{1,1}$ [7, A005314].
@@ 4. Questions
One way to extend our work is to study a similar game using a die instead of a coin. For example, if $E$ is the expected number of times to throw a die with $c$ faces marked $1,\ldots,c$ to obtain a run of $k$ ones, then $E=\sum_{i=1}^{k}\frac{(c-1)(E+i)}{c^{i}}+\frac{k}{c^{k}}$, and solving this equation with the help of the formula $\sum_{i=1}^{k}\frac{i}{c^{i}}=\frac{c}{(c-1)^{2}}-\frac{(c-1)k+c}{(c-1)^{2}\cdot c^{k}}$ yields $E=\frac{c^{k+1}-c}{c-1}$. Taking $c=2$ recovers our result $E(\mathrm{H}^{k})=2^{k+1}-2$ in Theorem 2.2.

### v2

@@ 1. Introduction
While all the above research work and other related online discussions occurred in 2024, Simonson discussed in a book published in 2022 [7, Ch. 2] a somewhat similar (but different) game in which a carnie offers \$5 to you while in a carnival and asks you to pay \$$n$ back, where $n$ is the number of times you flip a coin until you have two consecutive heads. Let $E$ be the expected value of coin flips to end this game. If the first flip gives a tail, then the expected value of flips to get two heads in a row will be $1+E$. If the first flip gives a heads and the next is a tails, then the expected value of flips will be $2+E$. If the first two flips both give a heads then the expected flips is $2$. It follows that $E=\frac{1}{2}(1+E)+\frac{1}{4}(2+E)+\frac{2}{4}$. This implies $E=6$, so in average you will lose \$1 each time when you play this game.
There is another way to compute $E$. Let $E_{n}$ be the number of outcomes from flipping a coin $n$ times such that the game will end right after $n$ flips. This turns out to be the Fibonacci number $F_{n-1}$ since $E_{n}=n-1$ for $n=1,2$ and for $n\geq 3$, if the first flip is a heads, then the next flip must be a tails, so by induction, $E_{n}=F_{n-3}+F_{n-2}=F_{n-1}$. It follows that the expected value of $n$ is $\sum_{n=1}^{\infty}nF_{n-1}/2^{n}$, which must equal $6$. This can also be proved by standard generating function technique (see, e.g., Vajda [9]), although Simonson did not address that, perhaps due to the scope of his book [7]. More generally, Benjamin, Neer, Otero, and Sellers [2] obtained some interesting Fibonacci sums by computing the $n$th power of the expected number of coin flips to obtain $\mathrm{H}\mathrm{H}$.
Simonson [7, Ch. 2] also mentioned that, since you will lose money eventually by playing the first game a number of times, what if you are offered to switch roles with the carnie but using HT instead of HH to end the game? Similarly as the above, one can show that the expected number of flips to end this game is $E=4$, so you should not get tricked again by the carnie. Interestingly, as mentioned by Simonson [7, Ch. 2], although the expected number of coin flips to get HH is more than the expected number of coin flips to get HT (and thus it is more likely for HT to outnumber HH in a long run), it is equally likely to get HT or HH when either occurs, since it is a $1/2$ chance to get either a heads or a tails after a heads.
@@ 3. Generalizations of Fibonacci numbers
Recall that the Fibonacci number of order $k$ is defined by $F_{n}^{k}=0$ for $n<k-1$, $F_{n}^{k}=1$ for $n=k-1$, and $F_{n}^{k}=F_{n-1}^{k}+F_{n-2}^{k}+\cdots+F_{n-k}^{k}$ for $n\geq k$. We have $F_{n,1}=1$ for $n\geq 0$, $F_{n,2}=F_{n}$ is the well-known Fibonacci number [8, A000045], $F_{n,3}$ is the tribonacci number [8, A000073], $F_{n,4}$ is the tetranacci number [8, A000078], and so on.
@@
Special cases include $\overline{F}_{n}^{1}=n+1$, $\overline{F}_{n}^{2}=F_{n+2}-1$ [8, A000071], $\overline{F}_{n}^{3}$ [8, A008937], $\overline{F}_{n}^{4}$ [8, A107066], and so on.
@@
For $m=0$ we have $F_{n}^{k,0}=2F_{n-1}^{k,0}-F_{n-k-1}^{k,0}$, so $F_{n}^{k,0}=\overline{F}_{n}^{k}$. The case $k=m=1$ [8, A005314] and the case $(k,m)=(2,1)$ [8, A059633] have also been studied before from other perspectives.
@@
We have $\widetilde{F}_{n}^{0,0}=F_{n+2}$, $\widetilde{F}_{n}^{k,0}=F_{n+1}^{k,1}$, and some other interesting special cases of $\widetilde{F}_{n}^{k,m}$, such as $(k,m)=(0,1)$ [8, A006053], $(k,m)=(0,2)$ [8, A158943], and $(k,m)=(1,1)$ [8, A112575].
@@
Special cases include $G_{n}^{2}=\widetilde{F}_{n-1}^{1,1}$ [8, A112575] and $G_{n}^{3}=F_{n-2}^{1,1}$ [8, A005314].
@@ 4. Questions
One of our summation identities (the first one in Corollary 3.3) partially overlaps with the sum $a_{n}=\sum_{k\geq 1}\frac{k^{n}F_{k}}{2^{k+1}}$ studied by Benjamin, Neer, Otero, and Sellers [2] similarly via both a probabilistic view and a generating function approach. It should be possible to generalize the above sum $a_{n}$ using higher order Fibonacci numbers or the two-parameter generalizations of Fibonacci numbers discussed in this paper.
Another way to extend our work is to study a similar game using a die instead of a coin. For example, if $E$ is the expected number of times to throw a die with $c$ faces marked $1,\ldots,c$ to obtain a run of $k$ ones, then $E=\sum_{i=1}^{k}\frac{(c-1)(E+i)}{c^{i}}+\frac{k}{c^{k}}$, and solving this equation with the help of the formula $\sum_{i=1}^{k}\frac{i}{c^{i}}=\frac{c}{(c-1)^{2}}-\frac{(c-1)k+c}{(c-1)^{2}\cdot c^{k}}$ yields $E=\frac{c^{k+1}-c}{c-1}$. Taking $c=2$ recovers our result $E(\mathrm{H}^{k})=2^{k+1}-2$ in Theorem 2.2.